# Jingjing Sun, Pengyi Shi, Jim Dai

### v3

@@ 6.5 Practical Implementation
For practical implementation, since our algorithm is model-based, the hospital first needs to calibrate the queueing system to reflect specific patient flows, arrival rates, and discharge patterns. Such calibration has been successfully applied in various hospital settings, e.g, Bertsimas and Pauphilet (2024). Once calibrated, managers can vary parameters such as holding and overflow costs, which act as levers to balance competing objectives: minimizing patient waiting or reducing overflow occurrences. By adjusting these values, managers can explore trade-offs and generate Pareto-optimal solutions to meet their desired operational goals. For instance, increasing overflow costs could reduce the amount of overflow, or vice versa.
@@
Note that the PPO algorithm employs a randomized policy that provides action probabilities for patient routing decisions. Decision-makers can use these probabilities, $\pi(f|s)$, to guide overflow strategies, such as selecting the most likely action or using de-randomization techniques (e.g., adding a final layer to the policy network for deterministic choices). The randomized policy also offers flexibility for managers to adjust routing decisions in real time. For example, while the algorithm suggests the broad direction such as which wards or departments require the most overflow, managers can refine these suggestions based on real-world factors like patient conditions or staff availability. Ultimately, decisions regarding which individual patients to overflow will depend on the judgment of hospital managers, hospitalists, and charge nurses; our algorithm supports, rather than replaces, the overflow decision-making.
@@ 7 Conclusion
This paper presents a scalable and efficient PPO algorithm for managing patient overflow in large hospital systems. Overflow management becomes increasingly complex as the number of hospital wards and patient classes increases, imposing challenges that traditional methods like ADP struggle to handle at scale. Our approach addresses this scalability issue by leveraging atomic actions and a partially-shared policy network. Additionally, our tailored designs for the hospital-specific queueing structures allow the training of our algorithm efficient and reduce the need for extensive simulation data, which overcomes a common limitation of reinforcement learning algorithms. The case study results demonstrate that our PPO algorithm matches or outperforms existing benchmarks in systems with up to twenty hospital wards.

### v4

@@ 6.5 Practical Implementation and Policy Interpretability
Practical implementation. Since our algorithm is model-based, the hospital first needs to calibrate the queueing system to reflect specific patient flows, arrival rates, and discharge patterns. Such calibration has been successfully applied in various hospital settings, e.g, Bertsimas and Pauphilet (2024). Once calibrated, managers can vary parameters such as holding and overflow costs, which act as levers to balance competing objectives: minimizing patient waiting or reducing overflow occurrences. By adjusting these values, managers can explore trade-offs and generate Pareto-optimal solutions to meet their desired operational goals. For instance, increasing overflow costs could reduce the amount of overflow, or vice versa; see, e.g., Appendix 12.3.
@@
Note that the PPO algorithm employs a randomized policy that provides action probabilities for patient routing decisions. Decision-makers can use these probabilities, $\pi(f\mid s)$, to guide overflow strategies, such as selecting the most likely action or using de-randomization techniques (e.g., adding a final layer to the policy network for deterministic choices). The randomized policy also offers flexibility for managers to adjust routing decisions in real time. For example, while the algorithm suggests the broad direction such as which wards or departments require the most overflow, managers can refine these suggestions based on real-world factors like patient conditions or staff availability. Ultimately, decisions regarding which individual patients to overflow will depend on the judgment of hospital managers, hospitalists, and charge nurses; our algorithm supports, rather than replaces, the overflow decision-making.
Policy Interpretability. Using visualization and explainable AI techniques, we analyze the policy generated by our PPO algorithm and extract actionable insights for hospital managers. First, the learned policy aligns with managerial intuition in several key areas: it overflows more patients when source wards are congested and/or destination wards are underutilized; it is less aggressive with overflow during the day, anticipating discharges, and more proactive in the evening when discharges are rare; and it dynamically balances patient preferences with system-wide congestion levels when selecting overflow destinations. Beyond these intuitive behaviors, the analysis also uncovers more subtle, non-obvious network-level effects. For example, the policy proactively reroutes patients to mitigate anticipated future congestion in interconnected wards; jointly adjusts overflow decisions across patient classes sharing downstream resources, accounting for the cascading “chain effects” of the network structure. These results demonstrate that the PPO policy not only aligns with operational intuition but also leverages system-level complexity to achieve performance gains that would be difficult to replicate via manual decision rules. Full visualizations and detailed analyses are provided in Appendix 13.
Practical and Algorithmic Insights. A key strength of our approach lies in its domain adaptation—leveraging problem-specific characteristics to develop more effective and efficient algorithms tailored in the hospital overflow context. By integrating a tailored policy network, customized basis functions, and a batching setup, our PPO algorithm requires significantly less simulation data than typical reinforcement learning or ADP-based approaches. This not only accelerates training but also makes the method practical in environments with limited data or computational resources. The case study highlights that domain-aware design choices are much more critical to algorithm performance than ad-hoc neural network tuning. Our findings offer a systematic path for adapting general-purpose RL frameworks to complex operational systems.
Beyond algorithmic contributions, our policy visualization analysis also provides actionable heuristics for hospital managers, for example, overflow intensity should vary by time of day, and decisions should depend on the relative congestion of both source and target units rather than follow rigid rules. While such principles can support better decision-making in the absence of sophisticated analytics, our study also highlights the limitations. Human-devised heuristics often fall short in capturing the anticipatory dynamics and interconnected network effects of large hospital systems. Scalable, data-driven tools like ours can systematically learn and adapt to these complexities, going beyond what ad hoc rules can achieve.
@@ 7 Conclusion
Overflow management becomes increasingly complex as the number of hospital wards and patient classes increases, imposing challenges that traditional methods like ADP struggle to handle at scale. The approach developed in this paper addresses the scalability issue by leveraging atomic actions and a partially-shared policy network. Additionally, our tailored designs for the hospital overflow context allow the training of our algorithm efficient and reduce the need for extensive simulation data, which overcomes a common limitation of reinforcement learning algorithms. The case study results demonstrate that our PPO algorithm matches or outperforms existing benchmarks in systems with up to twenty hospital wards.
The atomic action + PPO framework is particularly effective in settings with the following structural features: (i) combinatorial action spaces involving multi-customer assignment decisions to multiple service pools; (ii) intertemporal congestion effects, where current decisions impact future system state and cost through queue dynamics; (iii) periodic, long-run optimization with arrival and departure patterns follow strong time-of-day or day-of-week cycles, which requires policies to adapt dynamically over time while still optimizing long-run performance. While we apply the method to inpatient overflow management, similar structures arise in other healthcare and service systems such as dynamic nurse assignment (Sun et al. 2025), call centers or customer support systems, emergency responder dispatching, ride-hailing (Feng et al. 2021), and long-term care admissions (Dong et al. 2025). In these contexts, our approach offers a scalable alternative to full enumeration or problem-specific heuristics.